# A Configuration-First Framework for Reproducible, Low-Code Localization

### v1

@@
Abstract.
Machine learning is increasingly permeating radio based localization services. To keep results credible and comparable, everyday workflows should make rigorous experiment specification and exact repeatability the default, without blocking advanced experimentation. However, in practice, researchers face a three-way gap that could be filled by a framework that offers (i) low coding effort for end-to-end studies, (ii) reproducibility by default including versioned code/data/configurations, controlled randomness, isolated runs, recorded artifacts, and (iii) built-in extensibility so new models, metrics, and stages can be added with minimal integration effort. Existing tools rarely deliver all three for machine learning in general and localization workflows in particular.
@@ Abstract.
In this paper we introduce LOCALIZE, a low-code, configuration-first framework for radio localization in which 1) experiments are declared in human-readable configuration, 2) a workflow orchestrator runs standardized pipelines from data preparation to reporting, and 3) all artifacts, such as datasets, models, metrics, and reports are versioned. The pre-configured, versioned datasets reduce initial setup and boilerplate speeding up model development and evaluation. The design, with clear extension points, let experts add components without reworking the infrastructure. In a qualitative comparison and a head-to-head study against a plain Jupyter notebook baseline, we show that the framework reduces authoring effort while maintaining comparable runtime and memory behavior. Furthermore, using a Bluetooth low-energy dataset we show that scaling across training data ($1\times$ to $10\times$) keeps orchestration overheads bounded as data grows. Overall, the framework makes reproducible machine-learning-based localization experimentation practical, accessible, and extensible.
localization, machine learning, reproducibility, workflow orchestration, low-code
journal: TOSEMccs: Software and its engineering Application specific development environments
1. Introduction
Background and motivation. Location-based services (LBS) underpin navigation, logistics, social applications, entertainment, and healthcare, and are currently largely powered by Global Navigation Satellite Systems (GNSS, ’GPS’) (Huang and Gao, 2018). Although recent advances have improved the performance of these systems, they still have clear limitations, such as in high-precision applications, complex outdoor environments, and indoor settings (Werner, 2014). This motivates radio-based localization that can operate where GNSS degrades or is unavailable. Historically, localization relied on analytical models and signal geometry. With growing application demands and environmental complexity, localization research is shifting from analytical, model-based methods to data-driven machine learning approaches (Trevlakis et al., 2023) that learn mappings directly from data, fuse heterogeneous signals, and adapt to device and environment variation.
@@ 1. Introduction
Gaps in current practice. Moving from analytical to machine learning approaches has shown promise in many settings, but raises practical challenges for everyday research. The complexity and heterogeneity of ML pipelines raise the barrier to entry. GUI/low-code platforms reduce coding effort, but can constrain advanced experimentation and fine-grained control (Witten et al., 2025; Demsˇar et al., 2013). Conversely, code-centric workflows provide maximum flexibility but, absent dedicated tooling, place the burden of reproducibility, environment control, and consistent evaluation on the researcher (Pineau et al., 2021; Heil et al., 2021; Henderson et al., 2019; Bouthillier et al., 2021; Gundersen et al., 202). In many localization studies, the primary novelty is a new model or training method evaluated on public datasets (Trevlakis et al., 2023; Burghal et al., 2020). Onboarding data and recreating baselines becomes setup overhead rather than the research focus. The broader ML literature shows that small changes to preprocessing, hyperparameters, software versions, or random seeds can alter conclusions and hinder fair comparison (Pineau et al., 2021; Heil et al., 2021). Beyond statistical variance, exact repeatability is fragile: hidden global state, non-deterministic operations, data-order effects, and environment or hardware differences can change outputs even when seeds and settings are fixed (Bouthillier et al., 2021; Henderson et al., 2019; Pham, 2020). Such issues are easy to miss and hard to diagnose.
@@
This challenge is not unique to radio localization but represents a critical bottleneck in many scientific domains now adopting machine learning (Heil et al., 2021). Fields such as computational biology and materials science face a similar gap where the transition from traditional models to complex, data-driven pipelines strains existing research practices (Heil et al., 2021). The lack of standardized, reproducible, and extensible workflows leads to duplicated effort, fragile results, and a high barrier to entry for domain experts who are not also software engineers (Wilson et al., 2017). Therefore, the framework presented in this paper addresses a fundamental software engineering problem: how to build research-specific development environments that enforce scientific rigor by default without sacrificing the agility needed for rapid experimentation.
@@
Proposed approach and contributions. To enable the development of high-quality generic models, the wireless localization research community could benefit from a framework enabling: (i) reproducibility by default, (ii) ease of use (low-code with minimal boilerplate), and (iii) flexibility that is built in, so advanced experimentation can be conducted without fighting the tool-chain. While many platforms address ease of use or extensibility, few make reproducibility the default alongside both. This missing combination is the focus of our work. This paper introduces LOCALIZE  [^1] https://github.com/sensorlab/localize, a low-code, configuration-first framework that makes ML-based radio localization practical, accessible, and reproducible by reducing initial effort, enabling novice and expert use, and supporting expert extensibility while promoting sound experimental practices. The contributions of this paper presents are as follows:
@@
In the remainder of this paper, we review related work in Section 2, outline the problem and requirements in Section 3, describe the proposed framework in Section 4, present a reference implementation in Section 5, and evaluate it against established platforms in Section 7. Finally, Section 9 concludes the paper.
2. Related Work
@@
2.1. The Challenge of Reproducibility in Machine Learning
@@
A consistent lesson from the literature is that ML results can be sensitive to seemingly minor choices, so reproducibility requires attention to both variability and exact repeatability. Bouthillier et al. (Bouthillier et al., 2021) quantify sources of variability in benchmarks and show that initialization, data sampling and order, and hyperparameter settings all affect outcomes. They recommend practices such as repeated runs and clearer reporting to separate algorithmic gains from noise. Building on this empirical picture, Henderson et al. (Henderson et al., 2019) demonstrate that non-determinism in benchmark environments and implementation details can materially change results, arguing for stricter methodology and documentation. Collectively, these studies establish that achieving computational reproducibility requires meticulous control over the entire experimental pipeline.
@@
In response, subsequent work outlines reporting and release requirements. Heil et al. (Heil et al., 2021) introduce tiered levels of computational reproducibility (bronze/silver/gold) with concrete requirements covering code, data, captured environments, and automation. At the community level, Pineau et al. (Pineau et al., 2021) describe the NeurIPS reproducibility program, checklists, a code policy, and a community challenge, which raised expectations for transparency across submissions.
@@
Despite these advances, practical support in general ML platforms remains uneven. Gundersen et al. (Gundersen et al., 202) report that reproducing published results often requires non-trivial user effort. This suggests a clear need for research workflows that embed controls, documentation, and repeatable execution as default, integral features rather than optional, post-hoc additions.
2.2. General-Purpose Experimentation Platforms
@@
Several general-purpose platforms aim to simplify ML experimentation, but they typically fail to resolve the tension between usability, reproducibility, and extensibility, forcing researchers to make compromises. For example, graphical tools like WEKA (Witten et al., 2025) and visual programming environments like Orange3 (Demsˇar et al., 2013) excel at lowering the barrier to entry in machine learning research in different domains. They allow for rapid prototyping without extensive coding. However, they externalize the burden of reproducibility, as rigorously capturing an end-to-end experiment, including precise library versions and sources of non-determinism, requires manual effort outside the core workflow.
@@
Declarative, code-centric platforms like Ludwig (Molino et al., 2019) advance this goal by using configuration files to define deep learning experiments. This approach streamlines workflows and supports extensibility. While powerful for neural networks, it is not primarily designed for the classical machine learning pipelines common in localization studies and, similar to graphical tools, leaves environment and data versioning as an external concern.
@@
2.3. Other Related Efforts
Specialized tools also contribute valuable insights, typically focusing on specific domains or aspects of experimental management. Some efforts focus on automating workflows for a specific domain, such as CSI fingerprinting localization (Tong et al., 2021) or multivariate time-series forecasting (Ye et al., 2024), that package common data transforms and models. While these reduce engineering for their domains, they do not generalize readily across methods or guarantee environment capture and data versioning by default.
Libraries such as MLXP (Arbel and Zouaoui, 2024) focus on lightweight, code-centric management of configurations, runs, and results to reduce overhead and encourage replicable practice. However typically rely on external mechanisms for full environment capture or dataset and and provenance management versioning. Considering this, exact replay remains a user-assembled composition of tools.
As highlighted in related work, there is a recognized need for reproducibility, usability, and flexibility in research. Existing approaches make progress in specific areas—such as task, specific automation, task libraries, or experiment management, but they tend to address these aspects only in isolation. In contrast, our work is designed to integrate all three goals: reproducibility by default, ease of use, and flexibility for advanced experimentation.
3. Problem statement
This work aims to design and build a reproducible, low-code experimentation suite for radio localization without precluding advanced experimentation. Today, no widely used approach makes reproducibility the default while also keeping implementation effort low and preserving a clear path to customization. In Section 1, we set three goals: (i) reproducibility by default, (ii) ease of use (low-code with minimal boilerplate), and (iii) easy customization. Each goal carries concrete implications for how the system should behave.
@@ 3. Problem statement
Figure 1. Outline of the proposed architecture consisting of three components: configuration, artifacts, workflow orchestrator:
4. Proposed framework
To address the problem stated in Section 3, we propose a low-code framework LOCALIZE for radio localization that makes reproducible experimentation the default while preserving a clear path to advanced use. The design maps to the requirements in Section 3: R1: low-code operation, R2: reproducibility by default, R3: a common evaluation layer, R4: built-in extensibility, and R5: fast setup and iteration.
4.1. Design overview
@@
Figure 2. User workflow supported by the framework, separated by experience level. Novice users modify configurations, intermediate users adapt experiments or add datasets, expert users implement new stages or models.:
4.2. User types
@@
These usage modes interoperate by design: experts contribute reusable components, intermediates compose them into experiments, and novices explore configurations safely. Because execution is isolated and deterministic, and artifacts are versioned for every run, ensuring everything remains reproducible and auditable. This aligns with our requirements defined in Section 3, specifically R1, R2, R3, and R4.
5. Reference implementation
This section details the implementation details of the LOCALIZE framework defined in Section 4 and the design alternatives we considered.
5.1. Backbone: orchestration, versioning, and configuration
@@
5.2. Language and evnironment
We considered Python, R, Java, and MATLAB for the core runtime. Although MATLAB still appears in localization work (Kosek-Szott et al., 2025), recent studies and toolchains lean heavily on Python (Roy and Chowdhury, 2021; Raschka et al., 2020). We chose Python for its mature ML ecosystem (NumPy/Pandas, scikit-learn, TensorFlow/Keras) and smaller coding footprint. The framework is distributed as a standard Python repository. Users clone it and create the environment from an environment.yaml file. Environments are managed with Conda (or Mamba), and dependencies are pinned in the environment file so installations are consistent across machines.
5.3. Implementing the three-stage design
@@ 5.4. Reproducibility
5.5. Preconfigured datasets
The repository includes 5 preconfigured datasets and their corresponding dataset-specific stages (Prepare, Featurize, and Split with several different cross-validation strategies) and experiment templates: Lumos5G (Narayanan et al., 2020) and UMU datasets for cellular based localization (Asensio-Garriga et al., 2025), LOG-a-TEC dataset for Bluetooth low-energy based localization (Bertalanič et al., 2022), CTW 2019 (Arnold et al., 2019) and CTW 2020 (Gauger et al., 2020) datasets for Wifi based localization . For experimenting with any of these datasets, a user can copy a template, edit params.yaml, and run a study without writing glue code or configuration boilerplate.
6. Evaluation methodology
In this section we elaborate on the methodology to evaluate the proposed LOCALIZE framework introduced in Section 4 along the following four aspects: (i) cross-tool capabilities using a qualitative comparison, (ii) developer effort by comparing our framework with a plain Jupyter notebook under controlled edits, (iii) runtime and resource usage by pipeline stage, and (iv) scalability as dataset size increases.
6.1. Cross–platform qualitative comparison
As the first step in the evaluation of LOCALIZE, we identify the closest platforms that are strong candidates to enable similar functionality as the proposed framework and dive deeper to understand specific differences. We identify WEKA (Witten et al., 2025), Orange3 (Demsˇar et al., 2013), Ludwig (Molino et al., 2019), and Jupyter notebooks in JupyterLab (Granger and Pérez, 2021) as the closest relevant platforms.
Table 1. Requirement definitions (R1–R5) used in the qualitative comparison. Wording is broadened from Section 3 to facilitate a fairer cross-platform comparison

### v2

@@
Abstract
@@
As machine learning (ML) becomes the backbone of critical applications—ranging from logistics to healthcare it becomes even more important to keep the results credible and comparable. Everyday workflows should make rigorous experiment specification and exact repeatability the default, without blocking advanced experimentation. However, in practice, researchers face a three-way gap that could be filled by a framework that offers (i) low coding effort for end-to-end studies, (ii) reproducibility by default including versioned code/data/configurations, controlled randomness, isolated runs, recorded artifacts, and (iii) built-in extensibility so new models, metrics, and stages can be added with minimal integration effort. Existing tools rarely deliver all three for machine learning in general and localization workflows in particular. In this paper we introduce LOCALIZE, a low-code, configuration-first framework for radio localization in which 1) experiments are declared in human-readable configuration, 2) a workflow orchestrator runs standardized pipelines from data preparation to reporting, and 3) all artifacts, such as datasets, models, metrics, and reports are versioned. The pre-configured, versioned datasets reduce initial setup and boilerplate speeding up model development and evaluation. The design, with clear extension points, let experts add components without reworking the infrastructure. In a qualitative comparison and a head-to-head study against a plain Jupyter notebook baseline, we show that the framework reduces authoring effort while maintaining comparable runtime and memory behavior. Furthermore, using a Bluetooth low-energy dataset we show that scaling across training data ($1\times$ to $10\times$) keeps orchestration overheads bounded as data grows. Overall, the framework makes reproducible machine-learning-based localization experimentation practical, accessible, and extensible.
I Introduction
Background and motivation. The engineering of modern data-driven systems has undergone a paradigm shift. As machine learning (ML) becomes the backbone of critical applications—ranging from logistics to healthcare—the software engineering (SE) community faces a growing challenge: the reproducibility crisis in research software [Pineau2021, Heil2021]. While ML provides powerful mappings from raw data to actionable insights, the development process remains largely ad-hoc, characterized by fragile codebases where small changes in global state or environment lead to inconsistent results [bouthillier_accounting_2021]. This issue is particularly acute in radio-based localization powering the majority of location-based applications such as navigation, logistics, social applications, entertainment, and healthcare [huang2018location]. As the radio-based localization research and engineering is transitioning from rigid analytical models to ML powered ones [trevlakis_localization_2023], the research and development tools and pipelines that must evolve to support this evolution.
@@ I Introduction
Gaps in current practice. Moving from analytical to machine learning approaches has shown promise in many settings, but raises practical challenges for everyday research. The complexity and heterogeneity of ML pipelines raise the barrier to entry as shown in a recent exploration of AI system development [dolata2023making]. The study highlights that the stochastic nature of ML introduces a unique cognitive load on developers that traditional SE practices do not adequately address. GUI/low-code platforms reduce coding effort, but can constrain advanced experimentation and fine-grained control [weka_book, orange_pdf]. Conversely, code-centric workflows provide maximum flexibility but, absent dedicated tooling, place the burden of reproducibility, environment control, and consistent evaluation on the researcher [Pineau2021, Heil2021, henderson_deep_2018, bouthillier_accounting_2021, gundersen2021do]. In many localization studies, the primary novelty is a new model or training method evaluated on public datasets [trevlakis_localization_2023]. Onboarding data and recreating baselines becomes setup overhead rather than the research focus. The broader ML literature shows that small changes to preprocessing, hyperparameters, software versions, or random seeds can alter conclusions and hinder fair comparison [Pineau2021, Heil2021]. Beyond statistical variance, exact repeatability is fragile: hidden global state, non-deterministic operations, data-order effects, and environment or hardware differences can change outputs even when seeds and settings are fixed [bouthillier_accounting_2021, henderson_deep_2018, pham_problems_nodate, 10855627]. Such issues are easy to miss and hard to diagnose.
@@
This challenge is not unique to radio localization but represents a critical bottleneck in many scientific domains now adopting machine learning [Heil2021]. Fields such as computational biology and materials science face a similar gap where the transition from traditional models to complex, data-driven pipelines strains existing research practices [Heil2021]. The lack of standardized, reproducible, and extensible workflows leads to duplicated effort, fragile results, and a high barrier to entry for domain experts who are not also software engineers [wilson2017good]. Therefore, the framework presented in this paper addresses a fundamental software engineering problem: how to engineer research-specific development environments that enforce scientific rigor by default without sacrificing the agility needed for rapid experimentation.
@@
Proposed approach and contributions. To enable the development of high-quality generic models, the wireless localization research community could benefit from a well engineered framework enabling: (i) reproducibility by default, (ii) ease of use (low-code with minimal boilerplate), and (iii) flexibility that is built in, so advanced experimentation can be conducted without fighting the tool-chain. While many platforms address ease of use or extensibility, few make reproducibility the default alongside both. This missing combination is the focus of our work and follows the principles within engineering within boundaries [10883021]. This paper introduces LOCALIZE  [^1] https://github.com/sensorlab/localize, a low-code, configuration-first framework that makes ML-based radio localization practical, accessible, and reproducible by reducing initial effort, enabling novice and expert use, and supporting expert extensibility while promoting sound experimental practices. The contributions of this paper presents are as follows:
@@
In the remainder of this paper, we review related work in Section II, outline the problem and requirements in Section III, describe the proposed framework in Section IV, present a reference implementation in Section V, and evaluate it against established platforms in Section VII. Finally, Section IX concludes the paper.
II Related Work
@@
II-A The Challenge of Reproducibility in Machine Learning
@@
A consistent lesson from the literature is that ML results can be sensitive to seemingly minor choices, so reproducibility requires attention to both variability and exact repeatability. Bouthillier et al. [bouthillier_accounting_2021] quantify sources of variability in benchmarks and show that initialization, data sampling and order, and hyperparameter settings all affect outcomes. They recommend practices such as repeated runs and clearer reporting to separate algorithmic gains from noise. Building on this empirical picture, Henderson et al. [henderson_deep_2018] demonstrate that non-determinism in benchmark environments and implementation details can materially change results, arguing for stricter methodology and documentation. Furthermore, [10855627] recently formalized the concept of ’reproducibility smells’ defined as patterns in the development lifecycle, such as hidden global states and lack of automated artifact tracking, that lead to fragile and non-deterministic results. Collectively, these studies establish that achieving computational reproducibility requires meticulous control over the entire experimental pipeline.
@@
In response, subsequent work outlines reporting and release requirements. Heil et al. [Heil2021] introduce tiered levels of computational reproducibility (bronze/silver/gold) with concrete requirements covering code, data, captured environments, and automation. At the community level, Pineau et al. [Pineau2021] describe the NeurIPS reproducibility program, checklists, a code policy, and a community challenge, which raised expectations for transparency across submissions.
@@
Despite these advances, practical support in general ML platforms remains uneven. Gundersen et al. [gundersen2021do] report that reproducing published results often requires non-trivial user effort. This suggests a clear need for research workflows that embed controls, documentation, and repeatable execution as default, integral features rather than optional, post-hoc additions.
II-B General-Purpose Experimentation Platforms
@@
Several general-purpose platforms aim to simplify ML experimentation while only partly alleviating the cognitive load of AI development [dolata2023making]. They typically fail to resolve the tension between usability, reproducibility, and extensibility, forcing researchers to make compromises. For example, graphical tools like WEKA [weka_book] and visual programming environments like Orange3 [orange_pdf] excel at lowering the barrier to entry in machine learning research in different domains. They allow for rapid prototyping without extensive coding. However, they externalize the burden of reproducibility, as rigorously capturing an end-to-end experiment, including precise library versions and sources of non-determinism, requires manual effort outside the core workflow.
@@
Declarative, code-centric platforms like Ludwig [ludwig_pdf] advance this goal by using configuration files to define deep learning experiments. This approach streamlines workflows and supports extensibility. While powerful for neural networks, it is not primarily designed for the classical machine learning pipelines common in localization studies and, similar to graphical tools, leaves environment and data versioning as an external concern.
@@
II-C Other Related Efforts
Specialized tools also contribute valuable insights, typically focusing on specific domains or aspects of experimental management. Some efforts focus on automating workflows for a specific domain, such as CSI fingerprinting localization [Tong2021] or multivariate time-series forecasting [ye2024mvts], that package common data transforms and models. While these reduce engineering for their domains, they do not generalize readily across methods or guarantee environment capture and data versioning by default.
Libraries such as MLXP [arbel_mlxp_2024] focus on lightweight, code-centric management of configurations, runs, and results to reduce overhead and encourage replicable practice. However typically rely on external mechanisms for full environment capture or dataset and and provenance management versioning. Considering this, exact replay remains a user-assembled composition of tools.
As highlighted in related work, there is a recognized need for reproducibility, usability, and flexibility in research. Existing approaches make progress in specific areas—such as task, specific automation, task libraries, or experiment management, but they tend to address these aspects only in isolation. In contrast, our work is designed to integrate all three goals: reproducibility by default, ease of use, and flexibility for advanced experimentation. We position LOCALIZE not merely as a domain-specific tool, but as an implementation of the engineering within boundaries principle proposed by [10883021]. In this paradigm, when software involves non-deterministic or stochastic components like ML models, the framework must provide rigid boundaries—or guardrails—to ensure the integrity of the process. LOCALIZE adopts this by using a configuration-first approach that separates the ’intent’ of the experiment from the ’implementation’ of the logic, thereby providing a structured environment where scientific rigor is a byproduct of the software architecture itself.
III Problem statement
This work aims to design and build a reproducible, low-code experimentation suite for radio localization without precluding advanced experimentation. Today, no widely used approach makes reproducibility the default while also keeping implementation effort low and preserving a clear path to customization. In Section I, we set three goals: (i) reproducibility by default, (ii) ease of use (low-code with minimal boilerplate), and (iii) easy customization. Each goal carries concrete implications for how the system should behave.
@@ III Problem statement
Figure 1: Outline of the proposed architecture consisting of three components: configuration, artifacts, workflow orchestrator.
IV Proposed framework
To address the problem stated in Section III, we propose a low-code framework LOCALIZE for radio localization that makes reproducible experimentation the default while preserving a clear path to advanced use. The design maps to the requirements in Section III: R1: low-code operation, R2: reproducibility by default, R3: a common evaluation layer, R4: built-in extensibility, and R5: fast setup and iteration.
IV-A Design overview
@@
Figure 2: User workflow supported by the framework, separated by experience level. Novice users modify configurations, intermediate users adapt experiments or add datasets, expert users implement new stages or models.
IV-B User types
@@
These usage modes interoperate by design: experts contribute reusable components, intermediates compose them into experiments, and novices explore configurations safely. Because execution is isolated and deterministic, and artifacts are versioned for every run, ensuring everything remains reproducible and auditable. This aligns with our requirements defined in Section III, specifically R1, R2, R3, and R4.
V Reference implementation
This section details the implementation details of the LOCALIZE framework defined in Section IV and the design alternatives we considered.
V-A Backbone: orchestration, versioning, and configuration
@@
V-B Language and evnironment
We considered Python, R, Java, and MATLAB for the core runtime. Although MATLAB still appears in localization work [matlab_survey], recent studies and toolchains lean heavily on Python [Roy2021MLSurvey, python_survey]. We chose Python for its mature ML ecosystem (NumPy/Pandas, scikit-learn, TensorFlow/Keras) and smaller coding footprint. The framework is distributed as a standard Python repository. Users clone it and create the environment from an environment.yaml file. Environments are managed with Conda (or Mamba), and dependencies are pinned in the environment file so installations are consistent across machines.
V-C Implementing the three-stage design
@@ V-D Reproducibility
V-E Preconfigured datasets
The repository includes 5 preconfigured datasets and their corresponding dataset-specific stages (Prepare, Featurize, and Split with several different cross-validation strategies) and experiment templates: Lumos5G [narayanan_lumos5g_2020] and UMU datasets for cellular based localization [asensio-garriga_university_2025], LOG-a-TEC dataset for Bluetooth low-energy based localization [morano_outdoor_2022], CTW 2019 [arnold_novel_2019] and CTW 2020 [gauger_massive_2020] datasets for Wifi based localization . For experimenting with any of these datasets, a user can copy a template, edit params.yaml, and run a study without writing glue code or configuration boilerplate.
TABLE I: Requirement definitions (R1–R5) used in the qualitative comparison. Wording is broadened from Section III to facilitate a fairer cross-platform comparison